Evaluate the polynomial for x = -2.
2x^2-5x+5

Answer by ewatrrr(24785)


Hi
2x^2-5x+5 Evaluate the polynomial for x = -2
2*(-2)^2 -5(-2) + 5 = 8 + 10 + 5 = 23